OpenStudy (anonymous):

−5x + 8y = 18. a) Find the x-intercept b)find the y-intercept

OpenStudy (hoblos):

to find the x-intercept, set y=0 and calculate x for y=0 -5x =18 x=-18/5 (-18/5,0) to find the y-intercept, set x=0 and calculate y for x=0 8y=18 y=18/8=9/4 (0,9/4)

OpenStudy (anonymous):

for the y , i dont get it ?